# MIDLMedical Imaging with Deep Learning

### v1

@@ 2 Method and Experiments
We compare three different architectures: UNet, UNETR and ViM-UNet (our contribution), implemented in torch-em [Pape()]. The UNet has 4 levels and 64 initial features, which double each level. For UNETR we use the ViT of Segment Anything [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollár, and Girshick] and the same decoder as UNet, consisting of of two conv. layers per level and a transposed conv.; with input from the previous decoder and the ViT output. We chose this simple implementation over skip connections as in UNet and original UNETR, which connect respective levels in encoder and decoder. We found no negative impact of this design. For the ViM-UNet we use a ViM encoder with bidirectional SSM layers. Similar to ViT this model operates on patches; we use a patch shape of 16. The decoder design is identical to UNETR. For UNETR and ViM-UNet we compare encoders of different size; Base, Large, Huge for ViT and Tiny, Small for ViM.
@@ 3 Results and Discussion
We would like to express our gratitude to Sartorius AG for support in this research through the Quantitative Cell Analytics Initiative (QuCellAI). We also gratefully acknowledge the computing time granted by the Resource Allocation Board and provided on the supercomputer Lise and Emmy at NHR@ZIB and NHR@Göttingen as part of the HLRN infrastructure. The calculations for this research were conducted with computing resources under the project nim00007. We thank Mahdi Enayati for the valuable inputs on vision transformers in microscopy. We thank the authors of ViM [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang], Segment Anything [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollár, and Girshick], nnUNet [Isensee et al.(2021)Isensee, Jaeger, Kohl, Petersen, and Maier-Hein] and U-Mamba [Ma et al.(2024a)Ma, Li, and Wang] for making their code publicly available.

### v2

@@ 2 Method and Experiments
We compare three different architectures: UNet, UNETR and ViM-UNet (our contribution), implemented in torch-em [Pape()]. The UNet has 4 levels and 64 initial features, which double each level. For UNETR we use the ViT of Segment Anything [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollár, and Girshick] and the same decoder as UNet, consisting of two conv. layers per level and a transposed conv.; with input from the previous decoder and the ViT output. We chose this simple implementation over skip connections as in UNet and original UNETR, which connect respective levels in encoder and decoder. We found no negative impact of this design. For the ViM-UNet we use a ViM encoder with bidirectional SSM layers. Similar to ViT this model operates on patches; we use a patch shape of 16. The decoder design is identical to UNETR. For UNETR and ViM-UNet we compare encoders of different size; Base, Large, Huge for ViT and Tiny, Small for ViM.
@@ 3 Results and Discussion
The work of Anwai Archit was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - PA 4341/2-1. We would like to express our gratitude to Sartorius AG for support in this research through the Quantitative Cell Analytics Initiative (QuCellAI). We also gratefully acknowledge the computing time granted by the Resource Allocation Board and provided on the supercomputer Lise and Emmy at NHR@ZIB and NHR@Göttingen as part of the HLRN infrastructure. The calculations for this research were conducted with computing resources under the project nim00007. We thank Mahdi Enayati for the valuable inputs on vision transformers in microscopy. We thank the authors of ViM [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang], Segment Anything [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollár, and Girshick], nnUNet [Isensee et al.(2021)Isensee, Jaeger, Kohl, Petersen, and Maier-Hein] and U-Mamba [Ma et al.(2024a)Ma, Li, and Wang] for making their code publicly available.